The following amounts were reported on the December 31, 2019, balance sheet:

Cash $ 36,000
Accounts receivable 99,000
Common stock 180,000
Wages payable 22,500
Retained earnings 360,000
Land 90,000
Accounts payable 19,500
Bonds payable 528,000
Merchandise inventory 75,000
Buildings and equipment, net of
accumulated depreciation
810,000


Required:

  1. Calculate working capital at December 31, 2019.
  2. Calculate the current ratio at December 31, 2019. (Round your answer to 1 decimal place.)
  3. Calculate the acid-test ratio at December 31, 2019. (Round your answer to 1 decimal place.)
Answer #1

working capital = current assets - current liability

= (36,000 + 99000 + 75000) - (22,500+19500)

= 210000 - 42000 = $168000

current ratio = current assets / current liability

= 210000 / 42000 = 5

acid-test ratio = quick assets - current liability

= (36,000 + 99000 ) - (22,500+19500)

= 135000 / 42000

= 3.21